# One Battle After Another: Probing LLMs’ Limits on Multi-Turn Instruction Following with a Benchmark Evolving Framework

### v2

@@ Appendix D Performance on Different Constraints
Hardest Constraints: The most challenging group by a significant margin is EXT, where a large performance gap separates GPT-5 and Gemini-3.0-Pro from all other models. This highlights that while models can be prompted to include keywords, they are exceptionally poor at adhering to specific frequency counts. Following closely in difficulty are LEN and CS. These constraints all demand a form of global planning and state tracking over the fine-grained words and characters throughout generation. This suggests that while models are fluent producers of text, their ability to maintain adherence to fine-grained structural and quantitative rules remains a significant limitation.
@@ Appendix F Sensitivity to Different LLM Synthesizers
We analyze the performance of models on different subsets of samples generated by different synthesizers in Table 8. Note that these results are simultaneously influenced by the user styles randomly sampled for each instance. Nevertheless, we observe that Gemini-3.0-Pro achieves better performance on data synthesized by Gemini-2.5-Flash, whereas other models find it more challenging.
@@ Appendix G Performance trend with different styles
We prompted GPT-4.1 to classify the dataset into two categories, i.e., formal and informal, based on the linguistic style of user queries. The performance comparison is presented in Figure 7. GPT-5 remains relatively stable across different styles, whereas other LLMs exhibit varying performances. Gemini-3.0-Pro, MiniMax-M2, Grok-5-Fast, DeepSeek-V3.2, and Llama-4 favor a more formal linguistic style characterized by clear intentions. Conversely, the remaining models show a preference for informal styles, where user queries are typically more engaging.

### v3

@@ Appendix D Performance on Different Constraints
Hardest Constraints: The most challenging group by a significant margin is EXT, where a large performance gap separates GPT-5 and Gemini-3-Pro from all other models. This highlights that while models can be prompted to include keywords, they are exceptionally poor at adhering to specific frequency counts. Following closely in difficulty are LEN and CS. These constraints all demand a form of global planning and state tracking over the fine-grained words and characters throughout generation. This suggests that while models are fluent producers of text, their ability to maintain adherence to fine-grained structural and quantitative rules remains a significant limitation.
@@ Appendix F Sensitivity to Different LLM Synthesizers
We analyze the performance of models on different subsets of samples generated by different synthesizers in Table 8. Note that these results are simultaneously influenced by the user styles randomly sampled for each instance. Nevertheless, we observe that Gemini-3-Pro achieves better performance on data synthesized by Gemini-2.5-Flash, whereas other models find it more challenging.
@@ Appendix G Performance trend with different styles
We prompted GPT-4.1 to classify the dataset into two categories, i.e., formal and informal, based on the linguistic style of user queries. The performance comparison is presented in Figure 7. GPT-5 remains relatively stable across different styles, whereas other LLMs exhibit varying performances. Gemini-3-Pro, MiniMax-M2, Grok-5-Fast, DeepSeek-V3.2, and Llama-4 favor a more formal linguistic style characterized by clear intentions. Conversely, the remaining models show a preference for informal styles, where user queries are typically more engaging.